OpenStudy (anonymous):

A plane is in line with a 1200m landing strip. The pilot measures the angles of depression of the ends of the strip to be 15.8º and 11.5º. How high up is the plane AND how far is it horizontally from the nearer end of the strip?

OpenStudy (campbell_st):

well start by drawing a diagram and find the missing angles  now use the law of sines to find the value of x \[\frac{1200}{\sin(4.3)} = \frac{x}{\sin(11.5)}\] when you find x, you can then find the height of the plane and the distance from the airstrip using trig ratios... you'll have a hypotenuse, and an angle...in a right triangle. hope it helps